OpenStudy (anonymous):

a toy factory makes a red toy, a green toy, and a blue toy. in a lot of 2,5000 toys, 500 were red, 500 were green, and 500 were blue. the toy factory has a defect rate of 4%, regardless of the color. if a toy is chossen at random, what is the probability it is blue or red, and defective?

OpenStudy (anonymous):

wait there is 2,500 toys and 500 of each of the three colours? Shouldn't it be 1500?

OpenStudy (anonymous):

@brokenbells83

OpenStudy (anonymous):

not sure, i did the p(a) + p(b) method and got... 500/25000 * 500/2500= 1000/50000= 1/50... but that isnt right

OpenStudy (anonymous):

no im asking what is the correct number of toys, i think u made a typo

OpenStudy (anonymous):

thats the question, sorry if i did

OpenStudy (anonymous):

ooops i did the green ones are 1,500

OpenStudy (anonymous):

So we have 500 red + 500 blue = 1000 marbles of the choice of colours Hence the chance of picking out a red/blue marble is:\[\bf \frac{ 1000 }{ 2500 }=\frac{ 2 }{ 5 }\]Now the probability that any of these 1000 red/blue marbles is defected is:\[\bf \frac{ 4 }{ 1000 }=\frac{ 1 }{ 250 }\]Therefore the combined probability of having randomly picked a blue/red defected marble is:\[\bf \frac{ 2 }{ 5 } \times \frac{ 1 }{ 250 }=\frac{1}{625}\]

OpenStudy (anonymous):

@brokenbells83 Do you understand?

OpenStudy (anonymous):

yess i do thank you:) @genius12

OpenStudy (anonymous):

i dont understand why it is 4/1000 i think it must be 4/100